The product of two consecutive even integers x and y is 34 less than 17 times their sum. Find the two integers. (Assume x < y.)
Answer by mananth(16949)


y=x+2
xy =17(x+y)-34
plug value of y
x(x+2)=17(x+(x+2))-34
x^2+2x=17(2x+2)-34
x^2+2x=34x+34-34
x^2+2x-34x=0
x^2-32x=0
x(x-32)=0
x=0, OR x=32
32 & 34
...
m.ananth@hotmail.ca